# Meta-Learning Objectives for Preference Optimization

### v2

@@ 3.1 Meta Learning PO objectives
Similarly to (Alfano et al., 2024), we parameterize $g$, $\psi$ and $\phi^{-1}$ as a one layer neural network with 126 hidden units and non-negative kernels, where the activation functions are equally split among:
$x,\;(x)_{+}^{2},\;x^{3},\;(x)_{+}^{1/2},\;(x)_{+}^{1/3},\;\log((x)_{+}),$
$e^{x},\;\tanh(x),\;\log(\mathop{\mathrm{clip}}(x)/(1-\mathop{\mathrm{clip}}(x))),$
@@
To search for the best $g$, $\psi$ and $\phi^{-1}$ within this class, we employ the OpenAI-ES strategy. Denote by $\zeta$ the parameters of $g$, $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (10) when using the parametrized $\psi$ and $\phi^{-1}$. Lastly, let $F(\zeta)$ be the expected cumulative reward of $\pi^{\zeta}$, i.e. $F(\zeta)=\mathbb{E}_{\tau\sim(\mu,\pi^{\zeta},P)}r(\tau)$. We then use Adam (Kingma & Ba, 2015) to update the parameters $\zeta$ using the estimated gradient. In practice, to compute (12), we sample 128 values of $\epsilon$ to obtain 256 perturbed objective functions. We then train 256 agents with the perturbed objective functions on an a preference dataset. To measure the value of each agent, i.e. $\widehat{F}(\zeta^{\prime})$ for all perturbed $\zeta^{\prime}$, we sample 100 trajectories on the target environment for each agent, and take the average cumulative reward as estimate for the value of the agent. Refer to Appendix E for further discussion and details on the ES methodology.
In our experiments, we consider both the case where we fix $g=\log$ and learn $\psi$ and $\phi^{-1}$, and the case where we learn all three functions. We also perform the evolution on both the two-step and the one-step MPO classes.
Base Mixed Quality Noisy ($\varepsilon=0.1$) Noisy ($\varepsilon=0.2$) Noisy ($\varepsilon=0.3$)
RRHF 2789 $\pm$ 285 2245 $\pm$ 134 1730 $\pm$442 749 $\pm$ 498 330 $\pm$ 552
SLiC-HF 3255 $\pm$ 66 2478 $\pm$ 54 2329 $\pm$289 1964 $\pm$ 116 1135 $\pm$ 224
DPO 3528 $\pm$ 58 2766 $\pm$ 89 3082 $\pm$80 2530 $\pm$ 97 1519 $\pm$140
IPO 3618 $\pm$ 44 2937 $\pm$ 85 3162 $\pm$ 66 2392 $\pm$136 1133 $\pm$ 115
CPO 3450 $\pm$ 55 2322 $\pm$ 208 2967 $\pm$ 58 2427 $\pm$ 44 2000 $\pm$35
ORPO 3087 $\pm$ 322 2500 $\pm$ 71 2841 $\pm$ 50 2359 $\pm$ 43 1953 $\pm$ 37
R-DPO 2606 $\pm$ 65 2107 $\pm$ 40 2099 $\pm$ 50 1740 $\pm$ 36 1667 $\pm$ 24
SimPO 3683 $\pm$78 3117 $\pm$ 185 2314 $\pm$ 752 118 $\pm$ 715 -3828 $\pm$ 341
SFT 3287 $\pm$ 62 2344 $\pm$ 40 2733 $\pm$45 2345 $\pm$ 37 2049 $\pm$33
LPO 3774 $\pm$ 102 2841 $\pm$ 46 3617 $\pm$ 69 2705 $\pm$ 370 1569 $\pm$ 156
With $g=\log$
1S-MPO (mixed-quality) 3206 $\pm$ 330 3153 $\pm$ 274 1319 $\pm$ 714 -1625 $\pm$ 944 -3967 $\pm$ 382
1S-MPO (noisy, $\varepsilon=0.1$) 3789 $\pm$ 60 3210 $\pm$ 60 3813 $\pm$ 47 3280 $\pm$ 83 3279 $\pm$ 83
2S-MPO (mixed-quality) 3595 $\pm$ 57 2785 $\pm$ 78 3197 $\pm$ 58 2487 $\pm$ 110 1687 $\pm$ 58
2S-MPO (noisy, $\varepsilon=0.1$) 3551 $\pm$ 58 2784 $\pm$ 63 3190 $\pm$ 62 2552 $\pm$ 94 1569 $\pm$ 122
With parametrized $g$
1S-MPO (mixed-quality) 3560 $\pm$ 333 3627 $\pm$ 79 3371 $\pm$ 410 3230 $\pm$ 259 2681 $\pm$ 251
2S-MPO (mixed-quality) 3736 $\pm$ 51 3202 $\pm$ 64 3488 $\pm$ 73 2992 $\pm$ 87 2253 $\pm$ 125
1S-MPO (noisy, $\varepsilon=0.1$) 3861 $\pm$ 79 3075 $\pm$ 87 3724 $\pm$ 59 2845 $\pm$ 365 1771 $\pm$ 107
2S-MPO (noisy, $\varepsilon=0.1$) 3701 $\pm$ 52 3178 $\pm$ 59 3490 $\pm$ 95 2886 $\pm$ 127 2074 $\pm$ 136
1S-MPO (noisy, $\varepsilon=0.3$) 3931 $\pm$ 69 3244 $\pm$ 55 3834 $\pm$ 82 3735 $\pm$ 84 3417 $\pm$ 82
Table 2: Three Legged Ant (TLA) Adaptation. Performance of popular and discovered MPO algorithms on TLA, for various dataset settings. For each algorithm-dataset combination, we report the average value and standard error of 25 trained agents. For discovered MPO algorithms, we report performance of one-step MPO (1S-MPO) and two-step MPO (2S-MPO). For each discovered MPO algorithm, we specify on which setting it was discovered and report its performance across all settings (with fixed hyperparameters). We underline the highest (or two highest if their confidence interval overlaps) average performance among the human-designed algorithm and report in bold the overall highest, for each setting.
@@ 4 MuJoCo Experiments
Our first set of experiments is carried out on continuous RL tasks in MuJoCo. In particular, we show the performance of all the algorithms presented in Table 1 across several settings and we compare it with the performance of our discovered objectives. To maximize computational efficiency, all our MuJoCo experiments are implemented in JAX (Bradbury et al., 2018) using the brax (Freeman et al., 2021) and evosax (Lange, 2022) libraries. We report our hyper-parameters in Appendix I.
@@ 4.1 Tasks
The offline preference optimization task is defined as follows. We train one agent (the original agent) in the original Ant environment, achieving a reward of 6000, and another (the target agent) in the TLA environment, which achieves a reward of 3900. For comparison, the original agent achieves a reward of 1700 in the TLA environment. We then generate a preference dataset of 1280 rows, each with two trajectories of length 1000 starting from the same state. Each trajectory is generated by either the original or the target agent, depending on the current setting. A Bradley-Terry judge ranks each pair of trajectories and declares a winner, based on their true cumulative reward.
We also consider training a randomly initialized agent on the Hopper environment, using a preference dataset with 5120 pairs of trajectories. This setting addresses the case where a new behavior has to be learned completely from the preference dataset and no prior knowledge of the task is available. Due to space constraints, we report the results of the experiments for this task in Appendix H.
We consider three variations of the preference dataset, each meant to represent a common issue of real world data.
@@
  • Noisy dataset: same as the base dataset but each label given by the judge is flipped with probability $\varepsilon$.
  • Mixed-quality dataset: each trajectory in the dataset is generated by an agent selected at random between the original and the target one. The resulting dataset will consist of, approximately, $25\%$ comparisons between two trajectories from the target agent, $50\%$ comparisons between a trajectory of the target agent and one of the original agent, and $25\%$ comparisons between two trajectories of the original agent.
Figure 1: SimPO
Figure 2: Discovered 1S-MPO (shuffled)
Figure 3: ORPO
Figure 4: Discovered 1S-MPO (noisy)
Figure 5: Plots of the absolute value of the gradient of SimPO, ORPO, and of two discovered objectives, i.e. 1S-MPO (shuffled) and 1S-MPO (noisy, $\epsilon=0.3$). The dots are sampled datapoints from the training distribution, with lighter points representing more recent datapoints during the course of training. In the bottom right and corner of each plot, SimPO’s and ORPO’s gradients go to zero, while the gradient of our discovered loss functions keep increasing. The effect of this on training is evident observing the datapoints: in ORPO and SimPO, the likelihood of the rejected points is not minimized, leading to decreased performance.
@@ 4.2 Results
We provide the results of our experiments for TLA in  Table 2, which reports the performance of several PO algorithm and of our discovered objectives. We performed a hyperparameter search for each algorithm-dataset combination and only report the performance of the best hyperparameters. All algorithms are run for 12 epochs over the preference dataset, with the exception of DPO, IPO and R-DPO, which are run for 2 epochs after 10 epochs of SFT.
First of all, we notice that none of the human-designed algorithms manages to recover the performance of the target agent and that most of them experience a large drop in performance in the mixed-quality and noisy settings.
@@
The first group within Table 2 shows that SFT plays a key role in the performance across different settings. SimPO, which does not contain an SFT term nor an SFT step, is at the top of the leaderboard on the base and the mixed-quality setting but performs poorly on all noisy settings. IPO and DPO, which do not contain an SFT term but have an SFT step, are among the top performers on the base, mixed-quality and low noise settings. Their performance finally drops when the noise level reaches 0.3. Lastly, the algorithms that present an SFT term in their objectives, e.g. CPO, ORPO and, obviously, SFT, exhibit a subotpimal performance in the base and mixed-quality settings but are much more robust to noise than the other algorithms.
@@
$\pi^{\star}\in\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\lambda\log\pi_{\theta}(\tau_{w}|x)-\log\pi_{\theta}(\tau_{l}|x)\big{]},$
which we call Linear Preference Optimization (LPO). LPO corresponds to SLiC-HF with $\delta=+\infty$ and obtains better results than most of the existing algorithms in Table 2, confirming that it is important to design objectives that do not flatten when $\pi(\tau_{w})>\pi(\tau_{l})$.
@@
Table 2 also reports the performance of our discovered objectives, for both the case where we set the monotonic transformation $g$ to be the logarithmic function and where we parametrize and learn it. We learn a separate objective for each dataset setting and report the performance of each learned objective on all settings.
Differently from the human-designed algorithms, the discovered objectives recover the performance of the target agent in multiple instances and are more robust to the mixed-quality and noisy settings. We note that allowing the evolution procedure to learn $g$ leads to a better performance in all dataset settings. Additionally, we have that the objectives discovered within the two-step MPO class always have a lower performance than those within the one-step MPO class. This is probably due to the ability to modify the SFT term in the one-step MPO class, which is not present in the two-step class.
When exploring the one-step MPO class with $g=\log$, our discovery procedure always recovers a variation of CPO. That is, we obtain an objective that can be approximated as
$\pi^{\star}\in$ $\mathop{\mathrm{argmax}}_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big{[}\alpha\log(\pi(\tau_{w}))$
$+\lambda\log\sigma\left(\beta\log(\pi(\tau_{w}))-\beta\log(\pi(\tau_{l}))\right)\big{]},$
where the coefficients $\alpha$ and $\beta$ depend on the setting. In particular, we have a low value for $\alpha$ and a high value for $\beta$ in the noisy settings, while we observe the opposite in the mixed-quality setting. These results confirm the observations made on the hand-crafted objectives, whereby objectives with an SFT term are more robust to noise and objectives without are more robust to mixed-quality trajectories. When we search the two-step MPO class with $g=\log$, we recover DPO.
Figure 5 provides a visualization of the gradient of some of the objectives discovered when we parametrize and meta-learn $g$. In particular, we show the objectives discovered within the one-step MPO class on the mixed-quality and noisy ($\varepsilon=0.3$) settings. For comparison, we provide the same plots for the ORPO and SimPO objectives. The hand-crafted algorithms present a larger gradient when $\pi(\tau_{w})<\pi(\tau_{l})$ and a smaller one when $\pi(\tau_{w})>\pi(\tau_{l})$, that is, they induce large updates when the data-point contradicts the current behaviour of the agent, and small otherwise.
@@
For additional results and analysis for the Hopper task, we refer the reader to Appendix H.
WR $s_{\text{WR}}$ LCWR $s_{\text{LCWR}}$ avg length
SLiC-HF 16.77 1.32 13.18 0.28 3016
ORPO 24.47 1.52 19.25 0.29 1679
CPO 25.96 1.54 21.07 0.25 1623
CoLPO 31.94 1.65 24.01 0.21 2391
Table 3: Alpaca Eval LLM results. We report Win Rate (WR), standard error for Win Rate ($s_{\text{WR}}$), Length Controlled Win Rate (LCWR), standard error for LCWR ($s_{\text{LCWR}}$) and average length of responses. We trained gemma-7b on dpo-mix-7k for 3 epochs. We then compared our model to GPT-4 using Llama-3.3-70B-Instruct as judge. All algorithms use the same $\beta=0.5$ and the same learning rate of $55e-$ 5. In the case of CoLPO and SLiC-HF, we used the same cutoff rate at 5.

### v3

@@ 3.1 Meta Learning PO objectives
Similarly to Alfano et al. (2024), we parameterize $g$, $\psi$ and $\phi^{-1}$ as a one layer neural network with 126 hidden units and non-negative kernels, where the activation functions are equally split among:
$x,\;(x)_{+}^{2},\;x^{3},\;(x)_{+}^{1/2},\;(x)_{+}^{1/3},\;\log((x)_{+}),\;e^{x},\;\tanh(x),\;\log(\mathop{\mathrm{clip}}(x)/(1-\mathop{\mathrm{clip}}(x))),$
@@
To search for the best $g$, $\psi$ and $\phi^{-1}$ within this class, we employ the OpenAI-ES strategy. Denote by $\zeta$ the parameters of $g$, $\psi$ and $\phi^{-1}$ and by $\pi^{\zeta}$ the final policy obtained optimizing the objective in (11) when using the parametrized $\psi$ and $\phi^{-1}$. Lastly, let $F(\zeta)$ be the expected cumulative reward of $\pi^{\zeta}$, i.e. $F(\zeta)=\mathbb{E}_{\tau\sim(\mu,\pi^{\zeta},P)}r(\tau)$. We then use Adam (Kingma & Ba, 2015) to update the parameters $\zeta$ using the estimated gradient. In practice, to compute (19), we sample 128 values of $\epsilon$ to obtain 256 perturbed objective functions. We then train 256 agents with the perturbed objective functions on an a preference dataset. To measure the value of each agent, i.e. $\widehat{F}(\zeta^{\prime})$ for all perturbed $\zeta^{\prime}$, we sample 100 trajectories on the target environment for each agent, and take the average cumulative reward as estimate for the value of the agent. Refer to Appendix˜G for further discussion and details on the ES methodology.
We consider both the case where we fix $g=\log\sigma$ and learn $\psi$ and $\phi^{-1}$, and the case where we learn all three functions. We perform the evolution on both the two-step and one-step MPO classes.
Figure 1: Data generation for MuJoCo experiments. For each pair of trajectories, which share the same starting state, two agents are chosen based on the data generating strategy. The trajectories are then judged by a Bradley-Terry Judge and, in the noisy setting, the labels are flipped with probability $\varepsilon$.
@@ 4 MuJoCo Experiments
Our first set of experiments is carried out on continuous RL tasks in MuJoCo. In particular, we show the performance of all the algorithms presented in Table˜4 across several settings and we compare it with the performance of our discovered objectives. To maximize computational efficiency, all our MuJoCo experiments are implemented in JAX (Bradbury et al., 2018) using the brax (Freeman et al., 2021) and evosax (Lange, 2022) libraries. We provide an implementation of our methodology here and report hyper-parameters in Appendix˜K.
@@ 4.1 Tasks
The offline preference optimization task is defined as follows. We train one agent (the original agent) in the original Ant environment, achieving a reward of 6000, and another (the target agent) in the TLA environment, which achieves a reward of 3900. For comparison, the original agent achieves a reward of 1700 in the TLA environment. We then generate a preference dataset of 1280 rows, each with two trajectories of length 1000 starting from the same state. Each trajectory is generated by either the original or the target agent, depending on the current setting. A Bradley-Terry judge ranks each pair of trajectories and declares a winner, based on their true cumulative reward. We consider three variations of the preference dataset, each meant to represent a common issue of real world data.
@@
  • Noisy dataset: same as the base dataset but each chosen/rejected pair of labels given by the judge is flipped with probability $\varepsilon$.
  • Mixed-quality dataset: each trajectory in the dataset is generated by an agent selected at random between the original and the target one. The resulting dataset will consist of, approximately, $25\%$ comparisons between two trajectories from the target agent, $50\%$ comparisons between trajectories of different agents, and $25\%$ comparisons between two trajectories of the original agent.
We also consider training a randomly initialized agent on the Hopper environment. This setting addresses the case where a behavior has to be learned from the preference dataset and there is no prior knowledge of the task available. We report the results of the experiments for this task in Table˜6.
@@ 4.2 Results
We provide the results of our experiments for TLA in  Table˜1, which reports the performance of several PO algorithm and of our discovered objectives. We performed a hyperparameter search for each algorithm-dataset combination and only report the performance of the best hyperparameters. All algorithms are run for 12 epochs over the preference dataset, with the exception of DPO, IPO, SimPO and R-DPO, which are run for 2 epochs after 10 epochs of SFT. We provide an additional noisy setting ($\varepsilon=0.2$) and performance for other existing algorithms in Table˜5 in Appendix I.1.
We notice that none of the human-designed algorithms manages to recover the performance of the target agent and that most of them experience a drop in performance in mixed-quality and noisy settings.
@@
The first group within Table˜1 shows that SFT plays a key role in the performance across different settings. SimPO, which does not contain an SFT term nor an SFT step, is at the top of the leaderboard on the base and the mixed-quality setting but performs poorly on all noisy settings. IPO and DPO, which do not contain an SFT term but have an SFT step, are among the top performers on the base, mixed-quality and low noise settings. Their performance finally drops when the noise level reaches 0.3. Lastly, the algorithms that present an SFT term in their objectives, e.g. CPO, ORPO and, obviously, SFT, exhibit a subotpimal performance in the base and mixed-quality settings but are much more robust to noise than the other algorithms.
Table 1: Three Legged Ant (TLA). Performance of existing and discovered MPO algorithms on TLA, for various dataset settings. For each algorithm-dataset combination, we report the average value and standard error of 25 trained agents. For each discovered MPO algorithm, we specify on which setting it was discovered and report its performance across all settings (with fixed hyperparameters). We underline the highest (or two highest if their confidence interval overlaps) average performance among the human-designed algorithm and report in bold the overall highest, for each setting.
Base Mixed Quality Noisy ($\varepsilon=0.1$) Noisy ($\varepsilon=0.3$)
RRHF (Yuan et al., 2023) 2789 $\pm$ 285 2245 $\pm$ 134 1730 $\pm$442 330 $\pm$ 552
SLiC-HF (Zhao et al., 2023) 3255 $\pm$ 66 2478 $\pm$ 54 2329 $\pm$289 1135 $\pm$ 224
DPO (Rafailov et al., 2024) 3528 $\pm$ 58 2766 $\pm$ 89 3082 $\pm$80 1519 $\pm$140
IPO (Azar et al., 2024) 3618 $\pm$ 44 2937 $\pm$ 85 3162 $\pm$ 66 1133 $\pm$ 115
CPO (Xu et al., 2024a) 3450 $\pm$ 55 2322 $\pm$ 208 2967 $\pm$ 58 2000 $\pm$35
ORPO (Hong et al., 2024) 3087 $\pm$ 322 2500 $\pm$ 71 2841 $\pm$ 50 1953 $\pm$ 37
R-DPO (Park et al., 2024) 2606 $\pm$ 65 2107 $\pm$ 40 2099 $\pm$ 50 1667 $\pm$ 24
SimPO (Meng et al., 2024) 3683 $\pm$78 3117 $\pm$ 185 2314 $\pm$ 752 -3828 $\pm$ 341
SFT 3287 $\pm$ 62 2344 $\pm$ 40 2733 $\pm$45 2049 $\pm$33
Our algorithms $\downarrow$
LPO 3774 $\pm$ 102 2841 $\pm$ 46 3617 $\pm$ 69 1569 $\pm$ 156
With $g=\log\sigma$
1S-MPO (mixed-quality) 3206 $\pm$ 330 3153 $\pm$ 274 1319 $\pm$ 714 -3967 $\pm$ 382
1S-MPO (noisy, $\varepsilon=0.1$) 3789 $\pm$ 60 3210 $\pm$ 60 3813 $\pm$ 47 3279 $\pm$ 83
2S-MPO (mixed-quality) 3595 $\pm$ 57 2785 $\pm$ 78 3197 $\pm$ 58 1687 $\pm$ 58
2S-MPO (noisy, $\varepsilon=0.1$) 3551 $\pm$ 58 2784 $\pm$ 63 3190 $\pm$ 62 1569 $\pm$ 122
With parametrized $g$
1S-MPO (mixed-quality) 3560 $\pm$ 333 3627 $\pm$ 79 3371 $\pm$ 410 2681 $\pm$ 251
2S-MPO (mixed-quality) 3736 $\pm$ 51 3202 $\pm$ 64 3488 $\pm$ 73 2253 $\pm$ 125
1S-MPO (noisy, $\varepsilon=0.1$) 3861 $\pm$ 79 3075 $\pm$ 87 3724 $\pm$ 59 1771 $\pm$ 107
2S-MPO (noisy, $\varepsilon=0.1$) 3701 $\pm$ 52 3178 $\pm$ 59 3490 $\pm$ 95 2074 $\pm$ 136
1S-MPO (noisy, $\varepsilon=0.3$) 3931 $\pm$ 69 3244 $\pm$ 55 3834 $\pm$ 82 3417 $\pm$ 82
Temporally-aware
TA-MPO (1) 3577 $\pm$ 45 2730 $\pm$ 63 3106 $\pm$ 62 1971 $\pm$ 35
TA-MPO (2) 3625 $\pm$ 49 3088 $\pm$ 69 3443 $\pm$ 53 1988 $\pm$ 39
TA-MPO (3) 3352 $\pm$ 57 2256 $\pm$ 44 2725 $\pm$ 46 1923 $\pm$ 27
@@
$\pi^{\star}\in\mathop{\mathrm{argmax}}\nolimits_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big[\lambda\log\pi_{\theta}(\tau_{w}|x)-\log\pi_{\theta}(\tau_{l}|x)\big],$
which we call Linear Preference Optimization (LPO). LPO corresponds to SLiC-HF with $\delta=+\infty$ and obtains better results than most of the existing algorithms in Table˜1, confirming that it is important to design objectives that do not flatten when $\pi(\tau_{w})>\pi(\tau_{l})$.
@@
Table˜1 also reports the performance of our discovered objectives, for both the case where we set the monotonic transformation $g$ to be the logarithmic function and where we parametrize and learn it. We learn a separate objective for each dataset setting and report the performance of each learned objective on all settings. Differently from the human-designed algorithms, the discovered objectives recover the performance of the target agent in multiple instances and are more robust to the mixed-quality and noisy settings. We note that allowing the evolution procedure to learn $g$ leads to a better performance in all dataset settings. Additionally, we have that the objectives discovered within the two-step MPO class always have a lower performance than those within the one-step MPO class. This is probably due to the ability to modify the SFT term in the one-step MPO class, which is not present in the two-step class.
Figure 2: SimPO
Figure 3: ORPO
Figure 4: Discovered 1S-MPO (shuffled)
Figure 5: Discovered 1S-MPO (noisy)
Figure 6: Absolute value of the gradient of SimPO, ORPO, 1S-MPO (shuffled), and 1S-MPO (noisy, $\epsilon=0.3$). The dots are sampled datapoints from the training distribution.
When exploring the one-step MPO class with $g=\log\sigma$, our discovery procedure always recovers a variation of CPO. That is, we obtain an objective that can be approximated as
$\pi^{\star}\in\mathop{\mathrm{argmax}}\nolimits_{\pi}\mathbb{E}_{(s_{0},\tau_{w},\tau_{l})\sim\mathcal{D}}\big[\alpha\log(\pi(\tau_{w}))+\lambda\log\sigma\big(\beta\log(\pi(\tau_{w}))-\beta\log(\pi(\tau_{l}))\big)\big],$
where the coefficients $\alpha$ and $\beta$ depend on the setting. In particular, we have a low value for $\alpha$ and a high value for $\beta$ in the noisy settings, while we observe the opposite in the mixed-quality setting. These results confirm the observations made on the hand-crafted objectives, whereby objectives with an SFT term are more robust to noise and objectives without are more robust to mixed-quality trajectories. When we search the two-step MPO class with $g=\log\sigma$, we recover DPO.
Figure˜6 provides a visualization of the gradient of some of the objectives discovered when we parametrize and meta-learn $g$. In particular, we show the objectives discovered within the one-step MPO class on the mixed-quality and noisy ($\varepsilon=0.3$) settings. For comparison, we provide the same plots for the ORPO and SimPO objectives. The hand-crafted algorithms present a larger gradient when $\pi(\tau_{w})<\pi(\tau_{l})$ and a smaller one when $\pi(\tau_{w})>\pi(\tau_{l})$, that is, they induce large updates when the data-point contradicts the current behaviour of the agent, and small otherwise.
@@
4.3 Including temporal awareness
We build an algorithm within the TA-MPO family using the insights gained in the previous section. In particular, we keep the standard SFT loss, which was rediscovered in all our experiments, and use SimPO for the PO component of TA-MPO, given its high performance and its ability to continue optimizing the policy even when $\pi(\tau_{w})>\pi(\tau_{l})$. That is, we define the objective
$\pi^{\star}\in\mathop{\mathrm{argmax}}\nolimits_{\pi}\mathbb{E}_{\mathcal{D}}\big[(1-\alpha(t))\log\pi_{\theta}(\tau_{w}|x)+\alpha(t)\log\sigma(\beta(\log\pi_{\theta}(\tau_{w}|x)-\log\pi_{\theta}(\tau_{l}|x))-\gamma)\big].$ (13)
We try three expressions for $\alpha$, designed to put most of the weight on the SFT component of (13) at the start of training and switch to PO towards the end of training:
$(1):\;\alpha(t)=t;\quad(2):\;\alpha(t)=t^{2};\quad(3):\;\alpha(t)=\sigma(20(x-0.75)).$
Table˜1 shows promising results for temporally-aware PO algorithms, as the objective in (13) with the version (2) of $\alpha$ matches or surpasses all human-designed algorithms in all settings.
Table 2: AlpacaEval LLM results. We report win-rates and standard error (length controlled win-rates and standard error in parenthesis) for three combinations of base model and preference dataset. We report in bold font the highest winrate (length-controlled winrate) for each column.
gemma-7bhttps://huggingface.co/google/gemma-7b, dpo-mix-7khttps://huggingface.co/datasets/argilla/dpo-mix-7k gemma-7b, capybara-7k§§§https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized mistral-7bhttps://huggingface.co/mistralai/Mistral-7B-v0.3, dpo-mix-7k
[…]